OpenStudy (jakyfraze):

Using the formula C = 2πr, find the circumference of a circle with diameter of 28 inches. A. 56 in. B. 28 in. C. 88 in. D. 44 in.

OpenStudy (jakyfraze):

@Nnesha can you please explain this to me

Nnesha (nnesha):

they already gave u the formula so just plugin the value remember radius is half of diameter r= d/2

OpenStudy (anonymous):

28/2=14 C= 2 PI 14 = 87.96 C. 88

OpenStudy (anonymous):

c

Nnesha (nnesha):

\[r=\frac{ d }{ 2}\] d=diameter r= radius so replace d with 28 to find radius

OpenStudy (jakyfraze):

@Nnesha so the answer is A?

OpenStudy (anonymous):

The answer is C

Nnesha (nnesha):

how did you get 56 ?

Nnesha (nnesha):

d/2 = what ? replace d with 28

OpenStudy (jakyfraze):

so decide 28/2? 14

Nnesha (nnesha):

yes right now 2 pi times 14

Nnesha (nnesha):

\[C = 2\pi r\] so substitute for 14 then solve

OpenStudy (jakyfraze):

2 pi x 14?

Nnesha (nnesha):

yes right

OpenStudy (jakyfraze):

oh okay. pi= 3.14 x 2 = 6.28. 6.28 x 14 = 87. 92. ANSWER 88 :D right?

Nnesha (nnesha):

yes right

OpenStudy (jakyfraze):

yes!! thank you :)

OpenStudy (jakyfraze):

jazakallah khair

Nnesha (nnesha):

:=)